OpenStudy (anonymous):

Find y when 2^x = 32^y , 32^x = 16(^y+3) 1. y = 2/3 2. y = 11/21 3. y = 13/21 4. y = 4/7 5. y = 10/21

OpenStudy (tkhunny):

Really big hint: 16 = 2^4 and 32 = 2^5

OpenStudy (anonymous):

thanks